I need help with this problem please! Concord Corporation has the following information available for accruals for the year ended December 31, 2020. The company adFor each situation, prepare the adjusting entry required at December 31. (Round answers to 0 decimal places, e.g. 5,275. If nFor each situation, prepare the journal entry to record the subsequent cash transaction in 2018. (Round answers to 0 decimal
Answer #1
1 Dec 31,2020 Utilities expense 445
      Accounts Payable 445
2 Dec 31,2020 Salaries and wages expense 2160 =3780*4/7
      Salaries and wages payable 2160
3 Dec 31,2020 Interest expense 237 =47400*6%*1/12
      Interest payable 237
4 Dec 31,2020 Accounts Receivable 260
       Service Revenue 260
5 Dec 31,2020 Accounts Receivable 6200
       Rent Revenue 6200
Part-2
1 Jan 11,2018 Accounts Payable 445
      Cash 445
2 Jan 4,2018 Salaries and wages expense 1620 =3780*3/7
Salaries and wages payable 2160
      Cash 3780
3 Jan 1,2018 Interest payable 237
      Cash 237
4 Jan 4,2018 Cash 260
      Accounts Receivable 260
5 Jan 2,2018 Cash 6200
      Accounts Receivable 6200